*Present:* Bucko, Bruce, Mullet, Si, Polar, Odd Bloke, Fluff_eei, Tim and Sadiq
h4. Moving Bong Installs/Patches
Bruce stated that Bong usually hosts patches before a LAN to enable members to download/update in their own time. He asked what the plans were for the current systems to support this.
Tim said that Bong was not up but that we could in theory move patches from Bong to Backus if we had disk access (to Bong's disks).
After various discussions on the sizes of disks, the exec decided to use Backus as the patch server.
*Action* Fluff_eei to get patches moved to backus.
*Action* Tim to move Bong's disk to Backus
*Action* Fluff_eei to get a backus account.
h4. LAN Power
Sadiq reported that following conversations with Neil Williams, ITS/Estates are now considering what power CompSoc may need when they consider plans for the renovation of LIB 1 and 2.
h4. Promoting CompSoc
*Action* Fluff_eei to produce LAN poster.
*Action* Si to send Bucko copy of his sheet once it is finished.
Mullet reported that Bloomberg were now OK to sponsor the ProgComp and they may be able to design and supply posters for it.
*Action* Mullet to put Bloomberg in contact with Si regarding posters.
Polar reported that IBM has responded to emails, we now need to arrange a meeting.
*Action* Polar to arrange meeting with IBM
*Action* Polar to talk to IBM regarding LPI exams
*Action* Polar to talk to Warwick Skills and Lifelong learning regarding LPI
Polar reported Accenture had made no reply.
*Action* Polar to get prizes for LAN
h4. Tech Team Report
Tim reported that migration from Bermuda and Twilight should be complete by week 5 which should free up space in the cupboard for more servers.
Sadiq was now in-charge of backups from Molotov - they are being SCP'd across to Backus for now.
Tim also raised the issue of finding another backup mail-relay since Faux has removed postfix from his machine. This needed to be external to Warwick and not in the same data centre as PAT.
*Action* Bruce, Tim and Sadiq are to find a solution to the mail-relay problem.
Mullet also asked Tim to form a Gaming Team to support the LAN.
*Action* Sadiq to join gaming team.
Tim completed by stating that the migration from Molotov should be complete by Week 10 of term 3.
Fluff_eei asked where responses to the Gaming Survey should go. After much deliberation and debate it was found that gaming@ was not sensible since this forwarded copies to all members. Fluff_eei is to get responses sent to him.
*Action* Fluff_eei to produce a spreadsheet of results from gaming survey.
*Action* All to check previous weeks Action Points
*Action* Si to buy 2 PSUs
*Action* Alll to go through Bugzilla and remove completed tickets which relate to them.
Sadiq reported that he had made progress with clothing, results to come to next week's meeting if back by then.
Mullet raised that issue that Tim may not be here for the 2007-2008 academic year.
Tim replied that there was currently zero probability of him being back at Warwick for a Masters. Therefore the exec would need to consider a replace Technical Officer.
After much debate, various views were expressed as to whether the Society should have a Tech Team officer. Discussion was adjourned until a later meeting.
Next Meeting is 15:00 next week (Tuesday).